# Pragmatics in the Era of Large Language Models: A Survey on Datasets, Evaluation, Opportunities and Challenges

### v1

@@ 6.2 Fine-Grained Evaluation of Pragmatics
Despite the growing number of datasets and benchmarks designed to evaluate specific aspects of pragmatic abilities in NLP models, particularly LLMs, there remains a critical need for a comprehensive, high-level framework for fine-grained pragmatic evaluation. Below, we outline key gaps in current evaluation practices and propose recommendations for advancing fine-grained pragmatic evaluation frameworks.
Metrics for Pragmatic Alignment.
Current evaluation metrics, largely based on traditional NLP tasks, fail to capture the nuances of how well LLMs handle context and social norms. More dynamic and context-sensitive measures, including human-in-the-loop assessments, are needed to assess how well models align with pragmatic expectations. For instance, metrics from psycholinguistics and psychometrics Shu et al. (2024) are useful to assess the LLM responses.
In-depth Analysis of LLM Outputs.
In §3, we drew a diverse landscape of the NLP tasks that the datasets cover. However, most studies treat pragmatics as a classification or generation task, or one of the aspects in their task settings, without closely examining how models arrive at their responses. A more in-depth, qualitative approach—potentially incorporating human feedback and error analysis—could offer deeper insights into where models succeed or fail in pragmatic reasoning. This is particularly relevant for tasks requiring subtle social inferences, implicature resolution, or sensitivity to power dynamics in conversation. We need better measures for evaluating the thought competence in LLMs van Dijk et al. (2023).
Expanding Data Resources.
Unlike syntactic or semantic tasks, pragmatic evaluation relies on rich, contextually grounded datasets, which are difficult to construct at scale. While crowdsourcing and expert annotation are viable solutions, they are resource-intensive. A possible direction is to leverage LLMs themselves to generate controlled datasets, though augmented by human validation.
@@ 6.3 Pragmatics Aids LLM Alignment
The alignment of LLMs refers to their ability to generate responses that are not only factually correct but also contextually appropriate and socially coherent Shen et al. (2023). For instance, speech act theory Searle (1969) highlights that meaning emerges from interaction, not isolated sentences—aligning with the communicative goals of LLMs.
Pragmatics in Instruction and Alignment.
@@
LLMs rely on prompts to infer user intent, making pragmatics essential for refining instruction-following. Studies have used pragmatic frameworks to evaluate and adjust model responses (Ruis et al., 2023; Sravanthi et al., 2024; Yerukola et al., 2024, e.g.,), yet they lack robustness in pragmatic inference, failing to recognize indirect requests or subtle shifts in social meaning Lee and Daniel (2024). By systematically incorporating pragmatic constraints and examples, LLMs can better interpret ambiguous user inputs and produce responses that align with conversational expectations Vaduguru et al. (2024).
Human-Computer Interaction.
@@
Pragmatics can improve human-AI dialogue and collaboration by ensuring contextually appropriate and socially aware responses. However, chatbot alignment remains limited, especially in handling ambiguity and implicit meaning Martinenghi et al. (2024). Incorporating pragmatic reasoning can enhance user satisfaction by making interactions more intuitive and responsive to nuanced human intentions, contributing to the continuous advancement of language models Vaduguru et al. (2024).
Multi-Agent and Collaborative AI.
@@
Pragmatics is crucial in multi-agent systems, where agents must infer intent from indirect cues Li et al. (2024). While a single LLM can role-play different characters with a well-crafted prompt, communication degrades when multiple agents interact, due to LLMs’ lack of true belief-state reasoning Zhou et al. (2024). Addressing this requires models that better handle information asymmetry and implicit intent in multi-agent dialogue.
6.4 LLMs Enhance Pragmatics Research
@@
Finally, we look back to the pragmatics research. Pragmatics is a core part of linguistics, and experimental approaches have become more common over the past two decades Sauerland and Schumacher (2025). Recent advances in LLMs offer new opportunities to accelerate pragmatic research.
Stimulus and Material Design.
@@
Experimental pragmatics relies heavily on carefully designed stimuli to test specific hypotheses Schwarz (2017). LLMs can facilitate this by generating diverse and context-rich materials that mirror real-world communication. This can significantly streamline the experimental design process and enable more nuanced investigations into pragmatic phenomena.
Supporting Data Annotation.
@@
High-quality annotation is crucial for pragmatics research but is often costly and time-consuming. LLMs can support the annotation process by providing preliminary labels that human annotators can refine. This hybrid approach has shown promise in previous studies Tan et al. (2024) but requires careful implementation to ensure annotation quality and reliability Nasution and Onan (2024); Rønningstad et al. (2024). Combining LLM-assisted pre-annotation with thorough human review Wang et al. (2024) offers a more efficient yet accurate approach to data annotation.
Recommendations for Future Research in Linguistic Pragmatics.
@@
Leveraging LLMs’ potential in pragmatics research requires collaboration between computational linguists, cognitive scientists, and NLP practitioners. Researchers should explore integrating LLMs into experimental design, material generation, and task development to simulate complex conversational contexts. Rigorous human evaluation and qualitative studies are essential for reliability and interpretability.
@@ 7 Conclusion
This survey provides a comprehensive overview of existing resources for evaluating pragmatic understanding in NLP. By categorizing datasets according to the pragmatics phenomena they target and analyzing their task designs with data collection methods and how they are evaluated with the models, we highlight current trends and challenges in this field. Our findings underscore the need for more nuanced and contextually rich evaluation benchmarks, especially as LLMs continue to evolve. We hope to provide a valuable guide for researchers and practitioners advancing pragmatic reasoning in NLP, promoting systems with more human-like communication and interdisciplinary collaboration.
@@
Survey Sources.
Our survey predominantly encompasses literature within NLP, as the scope is to review the resources for evaluating pragmatics in NLP models. Therefore, we did not look extensively for resources in linguistic studies. This might result in an omission of potential resources or enlightening evaluation methods, which could bring insights to current data collection processes and evaluations. We advocate for future surveys in reviewing the resources for pragmatic language understanding in linguistics.
NLP Tasks.
Our survey is tailored towards tasks that feature applicability in NLP. As a result, intrinsic pragmatic tasks, such as discourse modeling are not discussed. Nevertheless, they could still be valuable to look into, especially in terms of gaining insights to current evaluation.
Multilinguality.
The scope of our survey is confined to literature published only in English. Although we endeavored to include works addressing multilinguality (albeit written in English), it is conceivable that pragmatic datasets created in other languages are not incorporated. This language constraint might limit the inclusivity of diverse pragmatic phenomenons and methodologies that are unique to a specific language.
@@
In this paper, we examine the evaluation of pragmatics in NLP, particularly in the context of LLMs. Since pragmatic abilities are inherently human, discussing these abilities in LLMs could raise concerns about anthropomorphism, as noted in prior work on LLM evaluation Ma et al. (2024a, b). It is important to clarify that our discussion of pragmatics does not imply human-like cognition or consciousness in LLMs. Instead, we advocate for human pragmatic features as benchmarks to evaluate LLMs with the goal of enhancing their design, ultimately making them more user-friendly and useful for human interaction.
@@ Acknowledgements
The authors acknowledge the use of ChatGPT exclusively to paraphrase and refine the text in the final manuscript. We thank the members of the MaiNLP lab from LMU Munich for their constructive feedback. YJL and BP are supported by ERC Consolidator Grant DIALECT (101043235).
@@ Appendix A Overview of Surveyed Works
To compile this survey, we conducted a comprehensive review of recent literature on evaluating pragmatics in NLP, with a specific focus on the datasets. We focused on identifying works that address these aspects, using the keywords “pragmatics” and “datasets” by primarily looking at papers publicized by 31.12.2024 at ACL Anthology [^1] https://aclanthology.org/, supplementing the initial search with manual qualitative additions. A total of 57 papers were surveyed. We only focus on the papers that address the pragmatics (the linguistic terms), rather than the literal meaning of the term “pragmatic” (meaning practical), which we filter out in the survey.

### v3

@@ 6.2 Fine-Grained Evaluation of Pragmatics
Despite the growing number of datasets and benchmarks designed to evaluate specific aspects of pragmatic abilities in NLP models, particularly LLMs, there remains a critical need for a comprehensive, high-level framework for fine-grained pragmatic evaluation. Below, we outline key gaps in current evaluation practices and propose recommendations for advancing fine-grained pragmatic evaluation.
Metrics for Pragmatic Alignment. Current evaluation metrics, largely based on traditional NLP tasks, fail to capture the nuances of how well LLMs handle context and social norms. More dynamic and context-sensitive measures, including human-in-the-loop assessments, are needed to assess how well models align with pragmatic expectations. For instance, metrics from psycholinguistics and psychometrics Shu et al. (2024) are useful to assess the LLM responses.
In-depth Analysis of LLM Outputs. In §3, we drew a diverse landscape of the NLP tasks that the datasets cover. However, most studies treat pragmatics as a classification or generation task, or one of the aspects in their task settings, without closely examining how models arrive at their responses. A more in-depth, qualitative approach—potentially incorporating human feedback and error analysis—could offer deeper insights into where models succeed or fail in pragmatic reasoning. This is particularly relevant for tasks requiring subtle social inferences, implicature resolution, or sensitivity to power dynamics in conversation. We need better measures for evaluating the thought competence in LLMs van Dijk et al. (2023).
Expanding Data Resources. Unlike syntactic or semantic tasks, pragmatic evaluation relies on rich, contextually grounded datasets, which are difficult to construct at scale. While crowdsourcing and expert annotation are viable solutions, they are resource-intensive. A possible direction is to leverage LLMs themselves to generate controlled datasets, though augmented by human validation Long et al. (2024). For instance, recent works Cong (2024); Srikanth et al. (2024) use GPT to generate synthetic stimuli, with validations from human evaluators.
@@ 6.3 Pragmatics Aids LLM Alignment
Recent research emphasizes incorporating linguistic insights into LLM development Opitz et al. (2025); Brunato (2025). We also advocate insights from pragmatics for LLM alignment. The alignment of LLMs refers to their ability to generate responses that are not only factually correct but also contextually appropriate and socially coherent Shen et al. (2023). For instance, speech act theory Searle (1969) highlights that meaning emerges from interaction, not isolated sentences—aligning with the communicative goals of LLMs.
@@
Pragmatics in Instruction and Alignment. LLMs rely on prompts to infer user intent, making pragmatics essential for refining instruction-following. Studies have used pragmatic frameworks to evaluate and adjust model responses (Ruis et al., 2023; Sravanthi et al., 2024; Yerukola et al., 2024, e.g.,), yet they lack robustness in pragmatic inference, failing to recognize indirect requests or subtle shifts in social meaning Lee and Daniel (2024). By systematically incorporating pragmatic constraints and examples, LLMs can better interpret ambiguous user inputs and produce responses that align with conversational expectations Vaduguru et al. (2024).
@@
Human-Computer Interaction. Pragmatics can improve human-AI dialogue and collaboration by ensuring contextually appropriate and socially aware responses. However, chatbot alignment remains limited, especially in handling ambiguity and implicit meaning Martinenghi et al. (2024). Incorporating pragmatic reasoning can enhance user satisfaction by making interactions more intuitive and responsive to nuanced human intentions, contributing to the continuous advancement of language models Vaduguru et al. (2024).
@@
Multi-Agent and Collaborative AI. Pragmatics is crucial in multi-agent systems, where agents must infer intent from indirect cues Li et al. (2024). While a single LLM can role-play different characters with a well-crafted prompt, communication degrades when multiple agents interact, due to LLMs’ lack of true belief-state reasoning Zhou et al. (2024). Addressing this requires models that better handle information asymmetry and implicit intent in multi-agent dialogue.
6.4 LLMs Enhance Research in Pragmatics
@@
Finally, we return to the research in linguistic pragmatics. Pragmatics is a core area of linguistics that has increasingly embraced experimental methods over the past two decades Sauerland and Schumacher (2025). Recent advances in LLMs offer new opportunities to advance pragmatic research.
@@
Stimulus and Material Design. Experimental pragmatics relies heavily on carefully designed stimuli to test specific hypotheses Schwarz (2017). The recently evolving LLMs might be able to facilitate this by generating diverse and context-rich materials that mirror real-world communication. This can significantly streamline the experimental design process for pragmatic phenomena.
@@
Supporting Data Annotation. High-quality annotation is crucial for pragmatic research but is often costly and time-consuming. LLMs can support the annotation process by providing preliminary labels that human annotators can refine. This hybrid approach has shown promise in previous studies Tan et al. (2024) but requires careful implementation to ensure annotation quality and reliability Nasution and Onan (2024); Rønningstad et al. (2024). Combining LLM-assisted pre-annotation with thorough human review Wang et al. (2024) offers a more reliable approach to data annotation.
@@
Recommendations for Future Research in Linguistic Pragmatics. Leveraging LLMs’ potential in pragmatic research requires collaboration among computational linguists, cognitive scientists, and NLP practitioners. Researchers should explore integrating LLMs into experimental design, material generation, and task development to simulate complex conversational contexts. Rigorous human evaluation and qualitative studies are essential for reliability and interpretability.
@@ 7 Conclusion
This survey provides a comprehensive overview of existing resources for evaluating pragmatic understanding in NLP. By categorizing datasets according to the pragmatic phenomena they target and analyzing their task designs with data collection methods and how they are evaluated with the models, we highlight current trends and challenges in this field. Our findings underscore the need for more nuanced and contextually rich evaluation benchmarks, especially as LLMs continue to evolve. We hope to provide a valuable guide for researchers and practitioners advancing pragmatic reasoning in NLP, promoting systems with more human-like communication and interdisciplinary collaboration.
@@
Survey Sources. Our survey predominantly encompasses literature within NLP, as the scope is to review the resources for evaluating pragmatics in NLP models. Therefore, we did not look extensively for resources in linguistic studies. This might result in an omission of potential resources or enlightening evaluation methods, which could bring insights to current data collection processes and evaluations. We advocate for future surveys in reviewing the resources for pragmatic language understanding in linguistics.
NLP Tasks. Our survey is tailored towards tasks that feature applicability in NLP. As a result, intrinsic pragmatic tasks, such as discourse modeling are not discussed. Nevertheless, they could still be valuable to look into, especially in terms of gaining insights to current evaluation.
Multilinguality. The scope of our survey is confined to literature published only in English. Although we endeavored to include works addressing multilinguality (albeit written in English), it is conceivable that pragmatic datasets created in other languages are not incorporated. This language constraint might limit the inclusivity of diverse pragmatic phenomenons and methodologies that are unique to a specific language.
Multimodality. While our paper covers a few data points that go beyond the textual-based features, such as the image captioning tasks (Tsvilodub and Franke, 2023, e.g.,), the majority of the data is textual-based. This gap has also been discussed in our §6.1. As we mainly focus on NLP venues, some resources could be not covered. Therefore, we encourage future work to explore pragmatic evaluation across diverse modalities.
@@
In this paper, we examine the evaluation of pragmatics in NLP, particularly in the context of LLMs. Since pragmatic abilities are inherently human, discussing these abilities in LLMs could raise concerns about anthropomorphism, as noted in prior work on LLM evaluation Ma et al. (2024a, b). It is important to clarify that our discussion of pragmatics does not imply human-like cognition or consciousness in LLMs. Instead, we advocate to use human pragmatic features to build benchmarks, aiming to enhance their design, and to enable user-friendly and effective human-machine interactions.
@@ Acknowledgements
The authors acknowledge the use of ChatGPT exclusively to paraphrase and refine the text in the final manuscript. We thank the members of the MaiNLP lab from LMU Munich for their constructive feedback. YL and KJ are supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), Project-ID 281511265 – SFB 1252 “Prominence in Language” in the project C06 at the University of Cologne. ZG is supported by the National Science Foundation via ARNI (The NSF AI Institute for Artificial and Natural Intelligence), under the Columbia 2025 Research Project (“Towards Safe, Robust, Interpretable Dialogue Agents for Democratized Medical Care”). YJL and BP are supported by ERC Consolidator Grant DIALECT (101043235).
@@ Appendix A Overview of Surveyed Works
To compile this survey, we conducted a comprehensive review of recent literature on evaluating pragmatics in NLP, with a specific focus on the datasets. We focused on identifying works that address these aspects, using the keywords “pragmatics” and “datasets” by primarily looking at papers publicized by 31.12.2024 at ACL Anthology [^1] https://aclanthology.org/.
This was chosen as the primary source since it is the main publication platform for the *CL community and contains the most relevant works on the evaluation of pragmatics in NLP models. After this initial search, we applied additional filtering criteria to refine the selection. Specifically, we focused on papers that
    introduced datasets explicitly designed for evaluating pragmatic phenomena in NLP,
    discussed methodologies for assessing pragmatic abilities of language models, or
    provided empirical evaluations of NLP models in related tasks to pragmatics.
In addition, we conducted manual qualitative inspections to ensure the inclusion of relevant works that might not have been captured due to variations in terminology; we also included a few recent papers from other sources out of the ACL Anthology based on our expertise in the field. This resulted in a final section of 58 papers in the current version.
Appendix B Pragmatic Phenomena and their Corresponding Tasks
In this section, we present the pragmatic phenomena and their corresponding tasks based on the survey works and show their mappings in Table 1. We noticed that there are no 100 percent one-to-one mappings between the pragmatic phenomena and the tasks, i.e., a phenomenon could be evaluated in either task in reality. This shows the possibility of diverse formats to evaluate the pragmatics of LLMs, and calls for future development of finer-grained tasks.
Task Type Context and Deixis Implicature and Presupposition Speech Acts and Intent Recognition Discourse and Coherence Social Pragmatics
Multiple-Choice Question Hu et al. (2023); Sravanthi et al. (2024); Park et al. (2024a, b) Hu et al. (2023); Sravanthi et al. (2024); Zheng et al. (2021); Kim et al. (2024) Durmus et al. (2019) Sap et al. (2019)
Question Answering Li et al. (2023a); Qi et al. (2023); Min et al. (2020); Srikanth et al. (2024); Kim et al. (2024); Miao et al. (2024); Sap et al. (2020) Qi et al. (2023); Louis et al. (2020); Srikanth et al. (2024); Damgaard et al. (2021); Müller and Plank (2024); Kim et al. (2024) Li et al. (2023a); Zhang et al. (2018); Rashkin et al. (2019) Miao et al. (2024) Yang et al. (2021); Sap et al. (2020); Zhang et al. (2018)
Natural Language Inference Westera et al. (2020) Jeretic et al. (2020); Halat and Atlamaz (2024); Nizamani et al. (2024); Koyano et al. (2022); Cong (2024); George and Mamidi (2020); Pedinotti et al. (2022) Welleck et al. (2019) Westera et al. (2020) Welleck et al. (2019)
Sentiment Analysis Ollagnier (2024) Yang et al. (2021) Ollagnier (2024); Yang et al. (2021); Buechel and Hahn (2017); Oraby et al. (2016)
Image Captioning Tsvilodub and Franke (2023); Bao et al. (2022); van Miltenburg et al. (2016)
Reference Games De Vries et al. (2017); Monroe et al. (2017, 2018); Takmaz et al. (2020, 2023); Shaikh et al. (2023) Shaikh et al. (2023); Khani et al. (2018); Greco et al. (2023) Shaikh et al. (2023)
Additional Tasks Sileo et al. (2022) Kameswari et al. (2020) Godfrey et al. (1992); Reinig et al. (2024) Pandia et al. (2021); Sadlier-Brown et al. (2024); Reinig et al. (2024)
Table 1: Mapping from Pragmatic Phenomena (columns) to NLP Task Types (rows).